Following is the information of a product of a firm:
   Selling price per unit = $60;
   Variable expenses per unit = $25;
   Breakeven point volume is 2,000 units

Fixed expenses per month = ?
------------


The following information exists for a firm:
Selling price per unit = $50          Contribution margin ratio = 30%
Fixed expenses per month = $30,000
Desired operating income = $18,000

The total revenues necessary to earn the desired operating income is = ??

{Please show calculations. I can't for the life of me figure this out! TIA!}
Answer #1

1.

Contribution margin per unit = Selling price per unit - Variable cost per unit

= 60 - 25 = 35

Breakeven point = Fixed expenses/Contribution margin per unit

Fixed expenses/35 = 2000

Fixed expenses = 2000*35

= 70,000

2.

Revenue necessary

= (Fixed expenses + Desired income)/Contribution margin ratio

= (30,000+18,000)/30%

= 48,000/30%

= 160,000